### v1

# Human Attention During Localization of Memory Bugs in C Programs

Abstract
This paper presents a study of human visual attention during localization of memory bugs in C. Human visual attention refers to the mechanical processes by which we selectively process and prioritize information. Visual attention is important to study because it is central to what information people (who are sighted) use to solve a particular problem. Meanwhile, memory bugs are among the most common types of bugs in C programs that manifest as a variety of program faults. In this paper, we study human visual attention while people attempt to locate memory bugs in code. We recruit 11 programmers to locate between one and six memory bugs in three C programs for 1.5-2 hours each. In total we collected observations of 17 hours of programmer effort. The bugs in our study cover memory leaks, overflows, and double frees, which are among the most common memory bugs. We analyze the task outcomes in terms of success rate and related factors, patterns of visual attention overall such as what lines and functions are read, and finally we explore differences of visual attention patterns during success versus failure cases.
Index Terms:
bug localization, memory bugs, human attention, visual attention, eye tracking
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
1 Introduction
This paper presents a study of human attention during localization of C memory bugs. Human visual attention refers to the mechanical processes by which we selectively process and prioritize information [1]. Human attention can be measured through eye tracking, interface logs, and activity surveys. It is important to study because beliefs about human visual attention influence decisions such as personnel allocation, user interface design, and even the architecture of machine neural networks [2, 3, 4, 5, 6, 7, 8]. Human attention has also long been academically interesting for “greater” philosophical reasons as a window into the working of our own minds [9].
Meanwhile, memory bugs are one of the most common programming errors in C programs [10, 11]. Memory bugs are behind aberrant software behavior such as excessive memory use, slowdowns, security vulnerabilities, and crashes. Some memory bug behaviors such as buffer overflows and segmentation faults are so common as to be tropes in C programming lore [12, 13]. Yet while often the cause of a memory bug is relatively small—a missing free(), an array length miscalculation, etc.—memory bugs can be difficult for programmers to find. Complex programs can have many pointers active at a time and memory leaks can be hard to observe as they accumulate slowly in long-running programs. This difficulty has made memory bugs a risk in C programs and a major research focus [10, 14].
Three types of memory bug dominate. Cotroneo et al. [15] point out that “use of uninitialized data, buffer overflows, and memory leaks are preponderant” as causes of bug behavior related to memory. Part of what makes these difficult to find is that they can lead to intermittent problems or challenges with vague non-functional requirements like responsiveness. For example it may be possible to use uninitialized data for a long time before it causes a problem, even if it is not strictly legal. Memory leaks may manifest as slow degradation in user experience rather than a diagnosable error message. Better understanding how people find these types of bugs could help address the lion’s share of challenges in localizing C memory bugs.
In this paper, we conduct an experiment in which programmers read C code and find memory bugs. We extracted eight bug reports from three large, popular C programs. Each bug is related to the three dominant memory bug types described in the previous paragraph or was a common bug type in the open-source C repositories. We recruited 11 programmers; in the study each programmer worked for 1.5 to 2 hours to find as many of the six presented bugs as possible. During the study we used eye tracking hardware to observe each programmer’s visual attention, while also tracking other interactions. We measured programmer performance such as success rate in finding bugs and time required. We also used a minimally-invasive attentional state survey for a self-assessment of attentiveness. Taken together these data form a view of visual attention during the bug localization tasks in C programs.
We found that: 1) people spend 75% of their fixations on just 25% of the functions they fixate on; 2) even unsuccessful participants fixate on the correct buggy line, but fail to recognize it; 3) successful bug locators have higher regression rates and average shorter distances between fixations; 4) successful bug locators revisit and/or spend more time on the buggy line, while unsuccessful participants do not return and/or spend less time on it.
2 Background and Related Work
This section discusses studies of human attention, studies of the process of bug localization, and the use of eye tracking in program comprehension research.
2.1 Studies of Human Attention
Eye tracking has been used extensively to study human attention. For example, computational models of eye movement during reading have been developed to predict how a person’s eyes will move over a body of text [16, 17]. These models highlight that eye movement across text can be predicted by textual features, which cause people to move from word to word, but also pause and reread text to ensure understanding. Such models have also been applied to non-reading tasks, such as visual search [18]. Collectively, the modeling efforts have revealed that there are systematic patterns of eye movements that people follow when reading or performing other visual tasks. Consequently, eye tracking has also been employed to examine when and why people deviate from those stereotypical patterns, thereby identifying the oculomotor correlates of mindlessness (or mind-wandering) during reading and other tasks. Specifically, multiple studies have found systematic differences in eye movements when people are reading attentively vs. inattentively [19, 20, 21, 22]. Specifically, when people are mind-wandering while reading, their eye movements are less sensitive to the content and complexity of the text, moving over the text unpredictably and unsystematically.
These phenomena have also been studied in other contexts, such as watching video lectures to simulate online coursework [23], listening to audiobooks [24], and viewing images of scenes [24, 25]. Faber et al. [24] have highlighted that inattentiveness is not always accompanied by the same patterns of eye movements across various tasks, as they require differences in the focality/dispersion of visual attention. Therefore, the task itself must be considered when examining such differences. The present study will thus seek to identify systematic oculomotor patterns when people are localizing bugs in code in an effort to understand how people do and do not detect bugs, and what their eyes are doing when that happens.
2.2 Eye tracking in Program Comprehension
Eye tracking technology has been a bedrock of program comprehension research for at least two decades and recent systematic surveys have led to recommended best practices in using this technology [26, 27]. Over time simultaneous hardware cost decreases and quality increases have expanded access to eye tracking equipment and the number of papers using it [28].
Program comprehension itself has a long history as research discovering how people read and understand source code [29, 30, 31, 32, 33]. Sometimes formalized as the “concept assignment problem” [34], the idea is that reading code is the process of interpreting the low-level implementation details in code as high-level concepts about program behavior. Eye tracking helps investigate the mechanical process of reading those low-level details. For example, Aschwanden and Crosby [35] observed how programmers’ eyes moved while examining a snippet of code featuring a loop and equations. Bednarik and Tukiainen [36] focused on scanpaths (i.e., sequences of eye movements) as participants tried to make sense of two small programs. Their findings revealed that programmers’ mental strategies sharpen over time, as indicated by fewer shifts in visual focus. Later, Rodeghero et al. [37] turned their attention to how developers write documentation, using this information to inform advances in automated summarization tools. Followup examination of these results extracted and categorized different eye tracking movement strategies [38, 39]. Abid et al. [40] expanded on Rodeghero’s study [37], incorporating real-world project contexts and a realistic IDE interface (better interfaces are possible with newer eye tracking devices and infrastructure). Similar findings have been observed in related areas such as reading UML diagrams, with some diagram formats leading to more thorough program comprehension than others [41, 42].
A few papers have sought to combine programmer visual attention with neural network-based models to predict human behavior or change machine behavior to be more like human behavior. For example, Karas et al. [43] compared eye movement pattern strategies when reading code for different tasks, then used these results to design novel neural network architectures to do those tasks [5]. Likewise Bansal et al. [44] modeled eye movements as scanpaths and augmented the attention mechanism in a neural network to improve code documentation generation [45]. The idea of combining human with AI attention is starting to take root outside of software engineering, with preliminary evidence showing performance improvements in various tasks such as image recognition and text generation [46, 47].
2.3 Studies of Bug Localization
Bug localization is the task of finding the area of code that causes a bug behavior [48, 49, 50]. Bug localization is a popular research target because understanding how people find bugs could lead to better automated tools to help them. Studies of bug localization have revealed strategies roughly falling into either static or dynamic approaches. Static strategies are those reading code itself while dynamic approaches execute that code. Dynamic approaches are well-studied and include tactics like program logging, assertions, breakpoints, and profiling [51]. Static approaches are less well-studied, though are a common starting point for programmers, especially when the code exists in a large codebase with which the programmer is not yet familiar. This situation is common when onboarding new staff, for example [52, 53]. This paper is primarily concerned with static approaches and specifically the mechanical process of attention during bug localization. Related papers include eye tracking studies such as by Bednarik and Tukiainen [54] who analyze the effect of experience level, finding that people with more experience read more areas of code but in less depth than novices. Research indicates that eye movements differ between debugging code and reading for other tasks [55]. One difference is that programmers search code more often in some tasks than others [56]. Given that differences exist in different tasks, this paper focuses on the narrow task of localization of memory bugs in C programs to avoid an assumption that movement patterns for general tasks generalize to this specific task.
3 Experimental Design
Our experiments center around observing what people pay attention to while trying to statically locate memory bugs in C programs. We study both micro- and macro-level attention. By “micro level” we mean mechanical details of attention such as fixation and regression counts. By “macro level” we mean what types of information or content are people reading and in what order. By “statically” we mean by just reading the code only, without executing it. By “locate bugs” we mean identify the line of code where the bug behavior originates. We define these terms in more detail throughout this section.
3.1 Research Questions
Our research objective is to understand the visual attention of programmers localizing bugs in C programs and whether these patterns differ between when participants successfully locate the bug versus cases when the bug is not identified. We ask the following Research Questions (RQs):
  RQ1
What are the outcomes of the task completion in terms of success rates, confidence, and other measures of human performance?
  RQ2
What patterns of visual attention do people follow during the bug localization tasks?
  RQ3
What patterns of visual attention are associated with successful versus failed task completion outcomes?
The rationale for RQ1 is to understand the degree to which people are able to complete the tasks as we present them in the experimental setting. An experiment with a diverse set of participant skill and task difficulty would generally lead to a diverse set of task outcomes. Therefore we start our analysis of attention by reporting the landscape of task outcomes. Next, our rationale behind RQ2 is to understand the visual attention that people follow during bug localization of C memory bugs. We seek to understand both micro- and macro-level attention. While there have been studies of attention for other tasks, we focus specifically on memory bugs in C programs. Finally, our rationale for RQ3 is to connect RQ1 and RQ2 results: we seek to understand if there are patterns of attention that are more or less likely to lead to success. These results are important because they could lead to educational or tool support interventions.
3.2 Procedure Overview
Our experimental procedure is a single-group cohort [57], “Quadrant II” study described by Stol and Fitzgerald [58]. What this means is that we focus on a single-group common characteristic (people proficient in C programming) in a controlled environment (as opposed to “Quadrant I” field studies, for example). We chose a single-group design because we do not seek to distinguish human factors such as age or gender, and we chose a controlled environment to reduce experimental variables. We 1) curate a list of bugs that is representative of typical memory bugs in C programs, 2) recruit programmers proficient in C to read bug reports and source code from that list, and 3) observe what the people do in a controlled environment.
The participant’s objective for each bug was to read the bug report, read the source code for the project containing the bug behavior, and find the line(s) that the participant believes are the cause of the bug. We did not ask the participant to actually fix the bug or run the program as our research goal is to understand the process of static bug localization.
We curated a pool of eight bugs and we recruited 11 people. We present each person a maximum of six bugs to find over 1.5 to 2 hours. Three bugs were the same for all participants and three rotated from our pool. To reduce experimental variables, we conducted the study one person at a time, with each person physically present in a quiet computer laboratory setting using the same hardware and software. Note that because we are primarily interested in patterns of attention, we avoided distracting participants from the computer. We chose a minimally-invasive design where an author of this paper was available for experimental setup at all times, but this person did not intervene or ask the participant questions in an e.g. think-aloud form.
3.3 Subject Bug Reports
We selected a total of eight bug reports from three C libraries/programs. Our goal is for these bug reports to be representative of typical memory bugs in C. For a definition of “typical” we turned to Cotroneo et al. [15] who studied characteristics of bugs and found that memory bugs are almost always one of three types: uninitialized data, buffer overflows, and memory leaks. We use bugs of the buffer overflow and memory leak type, but used double frees in lieu of uninitialized data due to their prevalence in our datasets. Bugs also have a range of difficulty and severity. At the same time we must balance diversity of bug selection with the practical reality of human studies, as has been pointed out in retrospective studies e.g. by Siegmund et al. [59]. Participants will learn about a program and seek to reuse that knowledge on other bugs, they will have different prior knowledge, have time limits, and will fatigue.
Table I below shows the three C libraries/programs we used. These programs have a range of total size and domain. Total size is important because finding bugs in larger programs may be more difficult. Domain is important because coding style may differ depending on domain [60]. We then selected between one and three bugs from the bug report repository for each program. We selected easy, medium, and hard difficulty bugs, where we assessed difficulty using comments in the bug report conversation, area of code that needed to be changed to fix the bug, and our own judgment. We also ensured that all bugs were double-frees, buffer overflows, or memory leaks. See Section 3.7 for a discussion of the risks and limitations of our choice of bugs. Also see our reproducibility package linked in Section 5.
For each bug, we utilized 1) the bug report, 2) follow-up communications among programmers and users, 3) the version of the code containing the bug, and 4) the patch fix showing lines changed to fix the bug.
[…]
OpenSSL 1.4m security 980
redis 250k database 734
sway 50k graphical 475
TABLE I: Statistics for C libraries from which we selected bugs.
3.4 Study Participants
We recruited 11 participants for this study. The participants had between 3.5 and 10 years experience programming. Five were graduate students and six were undergraduates. The undergraduates were recommended as strong C programmers by faculty, though these faculty were not authors of this study and this study was not connected to any coursework (per IRB requirements). We compensated participants at the fully-loaded local market rate of US\$100/hr.
3.5 Procedure Details
The details of our study procedure boil down to: 1) each participant visited the laboratory in person at a scheduled time, one at a time, 2) the participant sat at the eye tracking computer to become familiar with the settings and review informed consent documents, and 3) read and locate C memory bugs for a maximum of 120 minutes. We gave each participant a maximum of six bug reports to read. Some of the bug reports were the same for everyone, while we rotated others (as a balance between diversity of samples and measuring consistency of participant behavior). The first two bug reports were the same for all participants (see online appendix Section 5.1.2). The third, fourth, fifth, and sixth bug reports rotated from the pool of eight.
Figure 1: The interface presented to participants. Area A is the bug report, area B is the code, area C is the file explorer, and area D is a text chat window with GPT-4o for AI help.
An example image of the interface participants saw is in Figure 1. To reduce experimental variables and ensure accurate eye tracking, we did not allow the participants to navigate away from the interface; we instead made necessary information available within the interface. The interface itself is based on the Eclipse 4.33.0 IDE with iTrace 0.2.0 [61]. We set the IDE to have four areas, shown in Figure 1. Area A is information about the bug including the bug report. We showed only the portion of the report explaining the bug symptoms. We did not allow the participants to see followup in which the exact bug location may be revealed. Area A is also where participants wrote their answers.
Area B is a typical code editing window. Area C is a typical file navigation area. For simplicity we had each bug localization task as a separate folder; clicking each folder would load the bug information into area A and reveal the relevant program’s code. We ensured each participant worked on only one bug at a time, though we interrupted the participant if he or she had not located a bug within thirty minutes and moved the participant to the next bug.
We regulated access to outside resources via a chat window with GPT-4o using the RemainAI [62] Eclipse plugin (Area D in Figure 1). This window was necessary because: 1) we did not allow focus to be changed from the IDE interface, 2) we did not want participants to find the bug’s solution from a solved bug report online, 3) in practice AI help is quickly becoming an important part of the programming process so participants expect this support, and 4) the plugin allows us to capture queries that people type. We set the plugin so that it would not automatically have access to the code and we asked participants to ask general coding help questions instead of details about the program. This restriction was necessary because GPT-4o could have memorized the bug report information during training.
We created a minimally-invasive popup window to appear every five minutes. This window asked “In the few seconds before this screen appeared, what were you thinking about?” with options to answer:
  • I was focused on the task.
  • I was thinking about my performance on the task or how long it is taking.
  • I was thinking about things unrelated to the task (e.g., daydreaming).
  • I was distracted by sights or sounds in my environment.
  • My mind was blank.
The purpose of this window is to measure self-reported attention level. This method is used in cognitive psychology to generate momentary subjective assessments of attentional states [63, 64, 65]. Responses to these screens can be used to examine which signals, both in terms of overt behavioral responses and physiological indices, co-occur with and precede moments of attentiveness vs. inattentiveness. Similarly, in the present study, we sought to understand whether certain behaviors/eye tracking metrics would be associated with self-reported inattentiveness by participants.
When a participant was ready with an answer to a bug report, he or she answered “what” and “where” questions. To answer “where”, he or she entered the line number location into a specified area in the bug report window (Figure 1, area A). The participant could enter a range of lines if he or she could not decide or believed more than one line was involved. To answer “what”, the participant entered a brief description in English of what he or she believed the problem to be. We also asked the participant two questions. First, to rate his or her confidence in the answer on a scale of 1-5 where 1 is least and 5 is most confident. Second, to rate his or her perceived difficulty on a scale of 1-5 where 1 is least and 5 is most difficult.
3.6 Data Processing
We collected seven types of information:
  • Eye tracking data (Tobii Pro Fusion 120Hz, iTrace 0.2.0)
  • Screen video capture (1080p, 30 FPS)
  • Communication with AI help (GPT-4o)
  • Self-reported attention level (five minute intervals)
  • Self-reported bug difficulty and confidence
  • Bug cause answered by participant (“what”)
  • Bug location answered by participant (“where”)
We graded the participants’ answers on a scale of 1-5 where 1 was very inaccurate and 5 was highly accurate. Attempts without answers received 0. A score of highly-accurate meant that the reported location was identical to the one in the official patch fix or we determined to be another fix of equal or higher quality. To minimize bias, we applied the rubrics shown in Tables II and III to decide scores.
Score Bug Identification and Explanation
5 Correctly identifies and explains bug with details.
4 Correctly identifies bug and explains with some details. Given more time, would have provided more details. On the right path.
3 Partly correct in identifying bug and explaining, or somewhat vague. Possibly some conceptual or detail errors.
2 Mostly incorrect in identifying bug and explaining, or extremely vague with no details.
1 Missed basics of bug or completely incorrect with little to no details.
TABLE II: Rubric for grading “what” questions.
Score Bug Location and Thought Process
5 Correct line, backed up by correct thought process.
4 Given more time, could have narrowed down to the correct line, but was close (correct function or call stack).
3 Mostly correct understanding of problem, some misunderstandings, limited idea where to look.
2 Incorrect understanding of problem or not much beyond bug report, but provided a location that follows logic of incorrect understanding.
1 Incorrect understanding of problem and vague location, random guess.
TABLE III: Rubric for grading “where” questions.
To process and analyze the data, we used the “Map Tokens” and “Generate Fixations” tools in iTrace Toolkit V0.2.2 to extract fixations, lines, and tokens [66]. The inputs to the iTrace Toolkit are: iTrace-Core data, iTrace Eclipse plugin data, and srcML archive files [67]. The fixation settings for generating fixations were: Fixation Filter: IVT [68], and the default settings for IVT: Velocity Threshold: 50, Duration (ms): 80. The iTrace Toolkit outputs a database containing fixation and gaze information. Then, we reviewed the screen recordings for each task and noted the x and y coordinates of each area of the screen (Figure 1, areas A-D). These coordinates varied for each task because we allowed each participant to configure their screen. We build on scripts from  [69], and the inputs are: databases from iTrace Toolkit, the “where” and “what” scores, screen region coordinates, and the self-reported confidence and difficulty scores.
3.7 Threats to Validity
Our study has several threats to validity. One threat is that the selected bugs may not fully represent all C memory bugs. We mitigate this risk by choosing bugs as described in Section 3.3. A second threat is the varying detail in GitHub bug reports. To address this, we reviewed each report and made minor edits to ensure answering the “what” and “where” questions was non-trivial. Another threat is that all participants are graduate or undergraduate students, who may not fully represent industry professionals. To mitigate this, we selected only those with C experience and accepted only highly recommended undergraduates. Another threat is that some participants wore glasses and participants were not reminded to keep their heads still. However, the iTrace toolkit is designed to handle these concerns by using a conservative approach to identifying gazes and fixations.
To address eye tracker drift [27] and occasional disconnections, we recalibrated the eye tracker between tasks and after any interruptions. An additional threat is that screen recordings often froze mid-session, affecting our ability to retrieve screen region coordinates and AI usage data. However, participants rarely changed window layouts after initial setup, and we used eye tracking data to confirm AI usage based on the amount of time the participant looked at the AI window. Also, the eye tracker was not tracking during Participant 11’s ladybug task, so this task is not included in eye tracking metrics, but the “where” and “what” scores for this task are included. Finally, Participant 11 did not complete the firefly task or answer the “what” and “where” questions within 30 minutes. Since we cannot infer their understanding, they received a score of 0, and we do not include this data when comparing successful versus unsuccessful tasks.
4 Experimental Results
In this section we present our experimental results as answers to RQ1-3 and supporting evidence.
4.1 RQ1: Task Outcomes
Overall, we found around a fifth of the tasks were completed successfully in terms of finding the correct bug location and identifying the correct reason for the bug. About 21% of the graded accuracy for “where” questions were 5, shown in Figure 2(f), which corresponds to a finding of “correct” (Table III). Also, only around 21% of the graded accuracy for “what” questions were 5, shown in Figure 2(c), which corresponds to a finding of “correctly” identifying the bug (Table II). However, these numbers rise to around 34% if we include answers clearly on the right track, only limited by time (i.e. accuracy scores of 4 or higher). These results show the difficulty of bug localization given low overall success, yet still point to the tractability of these bugs by people, considering that over half were at least on the right track despite only reading the bugs for a few minutes (57% of “where” accuracy scores were 3 or higher). Still, a minority did not show evidence that the problem would have been solved even given much more time — for only around 15% of tasks was there complete failure to find the correct location (score of 0 or 1 for “where” accuracy) and 23% for describing the problem (score of 0 or 1 for “what” accuracy), including total guesses or bug descriptions demonstrating key misunderstandings.
Refer to caption Refer to caption Refer to caption
(a) “What” Confidence (b) “What” Difficulty (c) “What” Accuracy
Refer to caption Refer to caption Refer to caption
(d) “Where” Confidence (e) “Where” Difficulty (f) “Where” Accuracy
Figure 2: Histograms of confidence, difficulty, and accuracy scores from all participants and tasks. Confidence and difficulty are self-reported by participants. Accuracy is graded on a rubric by the authors of this paper. “What” refers to the bug cause. “Where” refers to the bug location in code.
What Where
Confidence Difficulty Accuracy Confidence Difficulty Accuracy
What Confidence 1.000 -0.744 0.065 0.156 -0.208 0.099
Difficulty -0.744 1.000 0.083 -0.203 0.261 -0.038
Accuracy 0.065 0.083 1.000 0.414 -0.162 0.927
Where Confidence 0.156 -0.203 0.414 1.000 -0.616 0.458
Difficulty -0.208 0.261 -0.162 -0.616 1.000 -0.209
Accuracy 0.099 -0.038 0.927 0.458 -0.209 1.000
TABLE IV: Pearson correlation coefficient matrix. Significant values (except diagonal) at $\alpha=0.050$ are in bold.
What Where
Program Issue Num Bug Type Codename Time Confidence Difficulty Accuracy Confidence Difficulty Accuracy
sway 6861 leak ladybug 1338 3.77 2.41 3.36 3.20 3.40 3.55
redis 8299 overflow stonefly 1424 3.14 3.41 2.73 3.23 2.95 3.14
sway 3577 double free hornet 1640 2.67 4.00 3.08 3.17 3.83 3.00
redis 9800 leak silverfish 900 3.50 3.50 1.50 3.50 3.00 1.50
redis 4527 overflow mantis 1284 3.00 4.00 2.20 2.00 4.25 2.20
sway 3300 double free spider 510 5.00 1.50 4.00 4.50 2.00 4.75
openssl 24892 leak firefly 1480 3.60 3.50 2.40 1.60 4.00 1.83
openssl 20278 double free weevil 930 1.75 4.00 3.88 3.75 3.25 3.50
TABLE V: Mean time required (in seconds) and mean scores per bug. Column “Issue Num” is the issue number in the project’s bug tracking system, included here for reproducibility. Column “Codename” is a name we used in our experiments and in this paper’s text for more memorable record-keeping.
Table IV shows Pearson correlations among all scores for all tasks and participants, with significant t-test results in bold. Notably the self-reported confidence and difficulty for “what” questions are not correlated to the rubric-graded accuracy for either question. In fact, there were around 13% of tasks for which participants self-reported high confidence (4 or 5) on “what” questions when graded accuracy scores were low (1 or 2). Still, confidence and difficulty scores are correlated, implying that participants are answering earnestly (random answers would tend to lead to uncorrelated responses). This observation may be due to the Dunning-Kruger effect, wherein people report high confidence in their knowledge despite weak performance [70].
Other correlations include: 1) “where” accuracy to “what” accuracy, which is not surprising considering that a correct rationale likely leads to a correct bug location, 2) “where” accuracy correlates to “where” confidence, implying that people are better judges of their performance of locating a bug, a finding supported by 3) “where” confidence is correlated with “what” accuracy, in that cases where people are confident in the location, they are more likely to give correct explanations (though, as noted above, not necessarily the other way around).
Overall participants reported being focused on the tasks at hand. In the attention question (prompted every five minutes), almost nine out of ten responses indicated attentiveness, as shown in Table VI. At a minimum, we interpret this finding to mean that we do not have evidence that participants were excessively distracted by the environment or other outside factors.
Option Percent
I was focused on the task. 87.68%
I was thinking about my performance on the task or how long it is taking. 9.85%
I was thinking about things unrelated to the task. 1.48%
I was distracted by sights or sounds in my environment. 0.99%
My mind was blank. 0.00%
TABLE VI: Self-Reported Attention Level Popup Options
TABLE VII: Gaze metrics for different bugs. Mean number of fixations for all bugs is 853 over 1188 seconds, or an average of 43 per minute.
Bug Fixation Count Regression Rate Fixation Duration (ms) Fixation Lines Fixation Functions
Mean Median Mean Median Mean Median Mean Median Mean Median
ladybug 700 747 57 61 465 453 153 151 29 27
stonefly 1064 1149 51 57 449 437 199 199 32 33
hornet 1313 1560 52 58 412 426 248 215 30 12
mantis 687 798 43 39 484 413 201 209 18 22
firefly 1007 986 47 42 410 426 300 372 50 57
silverfish 638 638 42 42 402 402 200 200 29 29
spider 376 376 60 60 453 453 71 71 8 8
weevil 1037 636 60 64 422 419 163 131 13 12
TABLE VIII: Descriptive statistics of the number of top (most fixated on) functions that capture 75% of all fixations on functions
Bug Functions in Top 75% Percent in Top 75%
Mean Min Max SD Mean Min Max SD
ladybug 5.90 2 18 4.53 20% 11% 38% 8%
stonefly 3.91 2 7 1.51 16% 7% 43% 10%
hornet 10.17 2 46 17.57 31% 15% 43% 10%
mantis 3.60 1 7 2.30 21% 17% 33% 7%
firefly 7.17 1 12 4.75 14% 9% 22% 5%
silverfish 7.50 2 13 7.78 33% 25% 40% 11%
spider 1.50 1 2 0.71 25% 17% 33% 12%
weevil 3.25 3 4 0.50 34% 16% 60% 21%
all bugs 5.54 1 46 6.99 22% 7% 60% 12%
Figure 3: Bar charts showing the distribution of fixations over functions read by participants locating the bug firefly. Blue indicates the top functions that capture 75% of the total fixations on functions. Most reading effort is concentrated on a few functions.
4.2 RQ2: Visual Attention Patterns
We summarize broad patterns of visual attention in Tables VII - IX and Figures 3 - 5. Table VII shows: Most fixation counts are comparable to the average of 853, with the exception of bug spider which is much lower, at 376. The most likely explanation is that this bug is also the easiest, as Table V shows spider with the highest confidence, lowest difficulty, and highest accuracy. Unsurprisingly, the bug with the highest mean fixation count (hornet) was among the lowest confidence and tied for highest difficulty. In all cases, the mean and median fixation counts and regression rates per bug are relatively similar, indicating that outliers are not markedly affecting the mean. The mean fixation duration in milliseconds was stable throughout, with no clear association with bug difficulty.
TABLE IX: Descriptive statistics of the number of top (most fixated on) lines that capture 75% of all fixations on lines of code
Bug Lines in Top 75% Percent in Top 75%
Mean Min Max SD Mean Min Max SD
ladybug 46.6 16 115 27.56 31% 19% 62% 13%
stonefly 49.82 19 109 26.37 29% 16% 68% 15%
hornet 96.50 52 207 60.41 37% 29% 62% 12%
mantis 71.80 33 123 36.51 37% 32% 44% 5%
firefly 107.50 11 164 65.05 33% 15% 40% 10%
silverfish 71.50 15 128 79.90 36% 36% 36% 1%
spider 22.50 10 35 17.68 30% 28% 33% 4%
weevil 42.50 29 60 15.46 29% 20% 35% 7%
all bugs 64.24 10 207 45.84 32% 15% 68% 11%
Figure 4: Bar charts showing the distribution of fixations over lines read by participants locating the bug firefly. Blue indicates the top lines that capture 75% of the total fixations on lines. Reading effort is distributed more widely than for functions.
A novel observation of this paper is that people read several functions outside of the one where the bug is located (a mean of 29), though visual attention is concentrated among relatively few. Table VIII shows that programmers spend 75% of their visual effort on 22% of the functions that they view. In rough practical terms, for a typical bug a typical person read 29 functions but spent 75% of their visual effort on 6 of them. Figure 3 shows a representative example of this pattern with the bug firefly: The x-axis shows one bar for each function viewed and the height of the bar on the y-axis is the number of fixations. Six study participants did this task. And while the number of functions viewed varied substantially (min 11, max 91), the pattern is consistent of concentrated visual attention. Put succinctly, around 75% of visual attention is paid to 25% of functions that are read.
We observe that the distribution of visual effort over lines is concentrated to a lesser degree than over functions. Consider Table IX. Around 32% of the lines read occupied 75% of the total number of fixations on code lines. The example in Figure 4 for bug firefly shows this pattern. While some participants focused on a small number of lines (e.g., p10), it was common to spread out one’s visual attention more, as indicated with the longer sections marked in Figures 3 and  4 in blue.
Around 60% of the participants’ fixations were spent on code on average, though this number was highly variable. The remaining fixations mostly spent reading the bug report details (a near-trivial amount of time was spent navigating the IDE). Recall from Figure 1 that participants had code, report, navigation, and AI areas visible and that they were not permitted to leave the IDE. Figure 5 shows the percent of visual attention (measured by fixations) each participant spent on code during each bug. We caution against drawing strong conclusions generalizing participant behavior in terms of region fixated on because we observe a range from 10% to 88% for some bugs. Some participants spend the vast majority of their effort reading code, while others spend that effort reading the bug report.
A final observation is that relatively little outside help was requested, as measured by the number of queries people asked in the AI help window and visual effort spent on this area. Table X shows only a handful of queries were made, with 5 of 11 participants not requesting AI help.
Figure 5: Percent fixations on code for each bug. Each point represents one participant’s percent of time looking at code versus other regions while working on a particular bug. Code corresponds to area B in Figure 1. Note not all participants worked on all bugs.
TABLE X: Participant AI Query Counts and Percentages
Participant AI Queries Percent Gazes at AI Window
p1 0 0.00%
p2 5 7.89%
p3 0 0.00%
p4 6 4.09%
p5 8 7.46%
p6 0 0.18%
p7 0 0.05%
p8 2 2.03%
p9 4 2.09%
p10 2 1.44%
p11 0 0.00%
4.3 RQ3: Attention and Task Outcome
Figure 6: Percent fixations on a “correct bug line” from lowest to highest “where accuracy” score for each participant-task. More fixations on correct locations are observed with higher accuracy, but correct lines were still read even at low accuracy.
Overall we observed differences in terms of regression rate and mean euclidean distance between fixations for higher versus lower performance on tasks. We did not observe a difference for fixation count or duration, or unique lines or methods read. Our findings are conceptually similar to reports from Software Engineering and other research areas that higher or lower performance on tasks can be associated with greater focus on certain thematically-relevant components [71, 72, 73, 74, 75]. This section discusses the nature of these differences in our study and key caveats.
First consider Figure 6 which shows the percent of fixations on a “correct bug line.” A “correct bug line” is a line in the code that would have been marked as a correct answer to the “where accuracy” question (see Section 3.6 for definitions). Each bar in Figure 6 is a task by a participant. The bar height is the percent of fixations during that task that the participant looked at a correct bug line. While the percentages themselves should not be over-interpreted, what we observe is that even during tasks with “where accuracy” of 2 or 3, participants often fixated on correct bug lines at least briefly. This means that participants failed to mark these lines as bug locations even though we have evidence that they read those locations.
Figure 7: As ‘Where Accuracy” increases, the mean distance between fixations tends to decrease, and the regression rate tends to increase.
As an initial point for understanding differences in attention based on task performance, we calculated six metrics of visual attention and split them based on high and low performance. We define “high” as “where accuracy” scores 4-5 and “low” as “where accuracy” scores 1-2.
TABLE XI: Statistical summary of eye metrics and accompanying boxplots. Low group means accuracy scores $\leq$2. High group means scores $\geq$4.
Metric Group Mean Min Max SD p
fixation count Low 830.7 45.0 2101.0 516.7 0.284
High 982.3 281.0 2411.0 624.8
fixation duration Low 416.0 302.4 502.4 52.5 0.161
High 442.9 333.2 596.4 73.7
regression rate Low 47.5 15.6 67.1 14.6 0.006
High 57.8 21.0 70.5 12.2
mean euclidean dist. Low 141.5 115.5 171.4 14.2 0.003
High 126.9 106.3 148.2 11.5
unique lines read Low 203.7 34.0 392.0 116.4 0.255
High 175.2 36.0 385.0 89.6
unique methods read Low 30.9 5.0 126.0 30.0 0.320
High 22.3 3.0 58.0 15.7
Figure 8: Distribution of bugs per participant grouped by “Where Accuracy.” Accuracy level 2 was most common, but participants were in all groups. Asterisk shows three rounded accuracy scores.
To preserve a conservative analysis, we use five metrics from earlier in this paper: fixation count and duration, regression rate, unique lines and methods read. One additional metric is Mean Euclidean Distance, the mean distance from the x,y coordinates of each fixation to the next fixation. Higher mean distance indicates more eye movement range. We performed a Mann-Whitney U test of low/high scores for each metric. We chose this test to be conservative in not assuming a normal distribution and to be robust to outliers. Table XI shows comparisons for each metric. We observe a statistically-significant difference in regression rate and mean euclidean distance. We observe differences in other metrics but these differences are not statistically-significant. Figure 7 shows regression rate and mean euclidean distance averaged for each “where accuracy” score. The general pattern is that regression rate is higher as “where accuracy” increases, while mean euclidean distance is lower.
Consider the examples in Figures 9-12 to illustrate potential differences in success versus failure cases. In these figures, each bar is a fixation on code (specifically area B indicated in Figure 1). Areas with zero-height bars indicate fixations on other parts of the screen or offscreen. The height of the bar is the duration of the fixation in milliseconds. Colors indicate “correct bug lines” (recall that only one correct bug line is necessary even if multiple are possible). Figure 9 shows a successful case on bug stonefly while Figure 10 shows a failure case. Likewise Figure 11 shows a successful case for bug weevil while Figure 12 shows a failure case.
For bug stonefly, Figure 9 shows the participant reading a correct bug location at line 97 of rdb.c around the seven minute time mark. The participant takes another eleven minutes to read other locations and conclude the session and write a correct answer to the problem. In contrast, Figure 10 shows a different participant reading several areas of code including correct bug lines, but for fewer fixations and less time per fixation – ultimately that participant provided an incorrect answer. A similar pattern is evident in Figure 11 where a participant had several fixations on line 1222 of http_client.c around the 18 minute mark, then continued to read other areas of code, before regressing to that location again and writing it as an answer at the end of the session. And the failure case in Figure 12 again shows fixations on the correct location early in the session but for less time, with the participant ultimately providing an incorrect answer.
A caveat to these findings is summarized in Figure 8: the low and high outcome groups are not balanced by participant because some participants performed less well overall, so some portion of our measurement could be of participant skill versus globally-relevant signs of success or failure. Thus we confine ourselves to conclusions that: 1) we observe higher regression rate and lower mean fixation distance among successful bug location attempts in our study, 2) success or failure is not necessarily explained by having seen a correct bug line but may be better explained by regression rate, and 3) participants did not necessarily report an answer immediately after reading a correct bug line and instead read and reread other areas of code. We view these findings as evidence that visual attention differs during success or failure as a motivation for future work.
Figure 9: Participant p7, success at bug stonefly.
Figure 10: Participant p3, failure at bug stonefly.
5 Discussion / Conclusion
This paper makes the following key contributions:
  1. 1.
    We perform a study of human visual attention during bug localization, specifically for memory bugs in C programs. To our knowledge, this study is the first of human visual attention during this particular problem.
  2. 2.
    We describe characteristics of visual attention such as the number of functions and lines to which attention is paid, highlighting specific areas of code needed to make a bug localization decision.
  3. 3.
    We find preliminary evidence pointing to differences in attention associated with task success or failure. This evidence opens the door for future work on potential positive interventions.
  4. 4.
    We release all data including raw eye tracking observations, post-processed token-level attention, etc., for future researchers.
This work has implications to other areas. Researchers in context-aware neural models and “agentic” AI may be informed by our findings of the extent of code context that humans need to solve localization of memory bugs (RQ2), as a guide to what inputs are necessary for the approaches they design. Researchers in accessibility tools for visual impairments may also find these findings useful to guide what data should be prioritized. Likewise, designers of educational tools may find clues in how to detect when interventions may be needed from our findings and data on attention differences in successful or failure cases. Finally, we present new knowledge of bug localization processes in C, which is of scholarly and public interest.
Figure 11: Participant p11, success at bug weevil.
Figure 12: Participant p5, failure at bug weevil.
To encourage reproducibility of our results and as a resource for other researchers, we release all data via the following online appendix:
6 Acknowledgments
This work is supported in part by the NSF grants NSF CCF-2100035 and CCF-2211428. Any opinions, findings, and conclusions expressed herein are the authors’ and do not necessarily reflect those of the sponsors.